CAVITY AND CHIRPED PULSE ROTATIONAL SPECTRUM OF THE LASER ABLATION SYNTHESIZED, OPEN-SHELL MOLECULE TIN MONOCHLORIDE, SnCl.^{a}

\underline{G. S. GRUBBS II}, DANIEL J. FROHMAN, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Hall-Atwater Laboratories, 52 Lawn Avenue, Middletown, CT 06459-0180; and S. A. COOKE, Department of Chemistry, University of North Texas, 1155 Union Circle # 305070, Denton, TX 76203-5017.

The use of laser ablation source-equipped chirped pulse and Balle-Flygare type cavity spectrometers have been utilized to accurately measure multiple isotopologues of the tin monochloride molecule in the X^{2}Π_{\frac{1}{2}} state. The molecule has been synthesized by ablating tin foil in the presence of 0.3% Cl_{2} in Ar. Rotational constants, nuclear electric quadrupole coupling constants, and magnetic hyperfine constants for the many isotopologues will be discussed. Although rotational analyses of this molecule have been previously performed^{b}, this is the first high-resolution, microwave study of SnCl.

aSupport from CHE-1011214

bN. Badowski, W. Zyrnicki and J. Borkowska, J. Phys. B: At. Mol. Phys. \underline{20} (1987), 5931-5937